OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

terenzreignz (terenzreignz):

To find the equation of the line, we need its slope, and a point on that line. We already have a point (6, -1) We need its slope... So what is the slope of this line?

OpenStudy (anonymous):

two lines are perpendicular to each other if m1=-1/m2...where m1 and m2 are slopes

OpenStudy (anonymous):

here slope of this line=-3...

OpenStudy (anonymous):

so slope of our line will be -1/-3=1/3

OpenStudy (anonymous):

now use point slope form

OpenStudy (anonymous):

y-y0=m(x-x0) ---->y-6=1/3(x+1)

OpenStudy (anonymous):

-x+3y=19

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

got it @AlmightySosa300 ?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

eq of given line is line 3x + y = 7. It can also be expressed as y = -3x+7 comparing it with y=mx+c we find slope m= -3 therefore the slope of the line perpendicular to it i.e. \[m=\frac{1}{3}\]since required line passes through the point (6, −1). threfore \[x_{1}=6 \] \[y_{1}=-1 \] eq of required line in slope point form is given as \[y-y_{1}=m(x-x_{1})\] substituting the respective values in the formula we find \[y-(-1)=\frac{1}{3}(x-6) \rightarrow y+1=\frac{1}{3}x-2\] i.e. \[ y=\frac{1}{3}x-3 \rightarrow 3y=x-9 \rightarrow -x+3y=-9 \rightarrow x-3y=9\] which is the required eq of the line

OpenStudy (anonymous):

y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17